# Brainformer: Modeling MRI Brain Functions to Machine Vision

### v1

@@ Abstract
“Perception is reality”. Human perception plays a vital role in forming beliefs and understanding reality. Exploring how the human brain works in the visual system facilitates bridging the gap between human visual perception and computer vision models. However, neuroscientists study the brain via Neuroimaging, i.e., Functional Magnetic Resonance Imaging (fMRI), to discover the brain’s functions. These approaches face interpretation challenges where fMRI data can be complex and require expertise. Therefore, neuroscientists make inferences about cognitive processes based on patterns of brain activities, which can lead to potential misinterpretation or limited functional understanding. In this work, we first present a simple yet effective Brainformer approach, a novel Transformer-based framework, to analyze the patterns of fMRI in the human perception system from the machine learning perspective. Secondly, we introduce a novel mechanism incorporating fMRI, which represents the human brain activities, as the supervision for the machine vision model. This work also introduces a novel perspective on transferring knowledge from human perception to neural networks. Through our experiments, we demonstrated that by leveraging fMRI information, the machine vision model can achieve potential results compared to the current State-of-the-art methods in various image recognition tasks.
@@ 1 Introduction
Figure 1: Given a pair of images and fMRI signals, Brainformer can explore the local patterns of fMRI signals from brain regions and discover how these regions interact with each other. Best view in color
Deep learning has recently experienced a surge in the development of foundational language models, e.g., BERT [13], GPT-2, GPT-3 [6], RoBERTa [33], T5[50], BART[29]. These models are typically trained on extensive datasets, often employing self-supervision at a large-scale web database. They have showcased their efficiency across a wide range of downstream tasks, utilizing strategies like zero-shot, few-shot, or transfer learning to extend the boundaries of large-scale modeling toward achieving human-level intelligence. Using the capabilities of text foundation models, a line of studies [64, 48, 23] has illustrated the feasibility of image-text foundation models. They encompass two encoders, i.e., an image encoder and a text encoder, jointly trained through contrastive learning. Apart from generating visual embeddings for tasks centered on vision alone, these image-text models can also generate textual embeddings that share the same latent space.
@@
From the success of text as supervision for the visual model, we ask what had been posed in the early days of artificial intelligence: What if human brain behaviors can serve as the guiding force for the machine vision models?. To answer this question, we found that Functional Magnetic Resonance Imaging (fMRI) is a neuroimaging technique that measures changes in blood flow and oxygenation in the brain, allowing researchers to observe brain activities in real-time. fMRI has provided valuable insights into various aspects of human cognition and neuroscience. For example, fMRI can help identify the specific regions of the brain that are active during various cognitive tasks. Studies have shown that different intelligence-related tasks activate specific brain areas, such as problem-solving or memory. In addition, human intelligence is not just about the activity of individual brain regions but also about how different brain regions communicate. fMRI can reveal functional connectivity patterns, highlighting networks involved in various cognitive processes.
@@
While text typically reflects the outcomes of the recognition process, fMRI captures the dynamics of the cognitive processes involved. Consider, for instance, an examiner tasked with describing the context of an image, such as “A man is sitting on the chair.” Initially, they focus on the man and then shift attention to the chair, determining the interactions between these objects, specifically, the act of sitting. Subsequently, they conclude the context in the text form. When encountering this description for the first time, it remains unclear which object the examiner attended to first or whether other smaller objects influenced their perception. Notably, this information can be encoded in fMRI signals. By analyzing specific brain regions or regions of interest, we can explore the whole recognition process of the human brain. This work provides a promising approach to further studies and endeavors to bridge the gap between human intelligence and deep neural networks. The contributions of this work can be summarized as follows.
@@
Contributions of this Work: This paper presents a novel Transformer-based framework, named Brainformer, to analyze and leverage fMRI signals as human brain activities to supervise the machine vision learning model. We present a simple yet effective way to encode fMRI features. Then, a novel Brain 3D Voxel Embedding approach is introduced to preserve spatial signal information. The proposed Brainformer explores the patterns within specific regions of interest in the human brain that signify the human attention process via the Multi-scale fMRI Transformer module. It can learn the correlation between these regions, helping to mimic the human recognition process. By incorporating these capabilities, Brainformer reveals mechanisms in the human recognition system. Second, a new Brain fMRI Guidance Loss function is presented to distill information from fMRI extracted using Brainformer. It serves as a guidance signal to enhance the performance capabilities of vision models. Finally, Brainformer is designed for self-supervised learning and trained as an end-to-end deep network. It consistently outperforms previous State-of-the-Art Self-supervised learning methods across standard benchmarks, i.e., object detection, instance segmentation, semantic segmentation, and brain response prediction.
@@ 2 Related Work
Figure 2: Brainformer utilizes fMRI signals from specific brain regions as input, extracting the local features representing patterns within each region. The TransformerBlock measures the correlation among these regions to emulate brain activities. This information is subsequently transferred to the vision model through Contrastive Loss and Brain fMRI Guidance Loss.
Vision Pretraining. Pretraining on the large-scale dataset such as ImageNet [12], Instagram [39], or JFT [65] has become a popular approach for many visual recognition problems, e.g., classification, localization, or segmentation, etc. Despite this method has great success, it is limited by the number of annotated training data that is costly to collect. Recently, self-supervised training has been proposed to address this problem. For example, BEiT [4] inspired by BERT [13] in natural language processing to make a masked image model. Similarly, SimMIM [62] and MAE are proposed to randomly remove the image patches and then use the encoder-decoder model to regress the masked pixel values. Nonetheless, these methods use the vision modality as the only input and are unsuitable for tasks involving images and text.
@@
Vision-Language Pretraining. Significant advancements have occurred in vision-language pertaining [49, 23, 66, 45, 60, 58, 30, 63, 67, 38, 59]. This approach aims to encode both visual and text information into a model. Initial work, e.g., LXMERT [55], UNITER [9], and VinVL [68], depended on pre-trained object detection modules like Fast(er) R-CNN [51] to extract visual representations of objects. Subsequent studies, such as ViLT [27] and VLMo [5], tried to integrate vision and language transformers and then train a multimodal transformer from scratch. Recent work has proposed Image-Text foundation models that bridge the gap between image and text understanding. It is worth mentioning that CLIP [48] and ALIGN [22] leverage dual-encoder architecture pre-trained on noise image-text pairs using contrastive learning. It has achieved robust image and text representation, suitable for cross-model alignment tasks and zero-shot image classification. Another line of research [46, 60, 58] explores encoder-decoder models trained with generative losses, showcasing impressive performance in vision-language benchmarks while maintaining strong visual encoder capabilities for classification tasks.
Decoding functional MRI. Decoding visual information from fMRI signals has been studied for a decade [18, 56, 24, 11, 17]. Most of these studies aimed to explore the hidden information inside the brain. It is a difficult task because of the low signal-to-noise ratio. Recently, with the help of deep learning, the authors in [10, 52, 54, 44, 31, 43] proposed methods to reconstruct what humans see from fMRI signal using diffusion models. There are limited works [42, 40] that incorporate human brain activities for transfer learning. However, these methods still need to be improved in exploring the working mechanism of the human brain.
@@
Different from recent studies on decoding fMRI [54, 52, 10], our goal is not only to explore valuable vision information from fMRI signal, i.e., semantic and structure information but also utilize them as supervision for helping to enhance the recognition capability of the vision model, a novel approach that standout our work from previous efforts in the field. Although our approach also aims to extract feature representations of fMRI data, these previous methods still have limitations and do not apply to our problem. Firstly, prior approaches, such as those by Yu et al. [54] and Paul et al., [52] relied on linear models and MLP layers to generate fMRI feature representations. This approach cannot capture the local pattern of the signals, especially the non-linear or complicated patterns. Kim et al. [26] presented SwiFT for self-training fMRI in 4D data, which does not apply to our problem. On the other hand, MindVis [10] utilized a Masked Brain Model inspired by MAE, which explored local patterns but ignored the correlations between multiple voxels, which describes how neurons interact inside the brain. Furthermore, MindVis failed to leverage the 3D spatial information in fMRI signals, which is crucial in understanding the relationships between voxels. Most notably, none of these approaches leverage characteristics of functional regions of interest (ROI) within the brain, which contain rich information about visual stimuli.
@@ 3 The Proposed Approach
3.1 Motivations
The fMRI contains a structured and semantically rich representation of information since fMRI captures brain activities in the context of a person engaging in visual tasks. It provides ground truth data about how the human brain responds to visual stimuli. According to recent studies [47, 25, 15], the brain can be divided into several Regions of Interest (ROI), where each region holds a different function. Especially in the scope of visual cognition, these are 6 regions specified as Early retinotopic visual regions (prf-visualrois), Body-selective regions (floc-bodies), Face-selective regions (floc-faces), Place-selective regions (floc-places), Word-selective regions (floc-words) and Anatomical streams (streams). In summary, these regions are responses for processing information related to the human body, face, place, motion, or objects regarding identification, recognition, etc. Therefore, the fMRI signals from these regions can be used to train computer vision models, providing a reference for what the brain is “seeing” during the training. By training vision models jointly with brain activities observed in fMRI scans, they can explore the relationship between the visual information processed by the brain and the visual information that models are learning. It can help uncover how the brain represents and processes visual information.
@@ 3.2.1 High-dimensional fMRI Feature Encoding
$\small\begin{split}\mathbf{r}^{k}&=\texttt{Conv1D}(m^{k})\in\mathbb{R}^{N^{k}\times d_{r}}\end{split}$ (2)
@@ 3.2.2 Brain 3D Voxel Embedding
$\small\begin{split}v_{i}^{k}&=\texttt{Linear}(x_{i}^{k},y_{i}^{k},z_{i}^{k})\in\mathbb{R}^{d_{r}}\\ v(m^{k})&=\texttt{concat}\left[v_{0}^{k},v_{1}^{k},\dots,v_{N^{k}-1}^{k}\right]\in\mathbb{R}^{N^{k}\times d_{r}}\end{split}$ (3)
@@ 3.2.3 Mutli-scale fMRI Transformer
  • We first employ a slicing window with $w$ of width that traverses the entire signal $\mathbf{r}^{k}$ from its beginning to the end, with a step size of $s$. This process decomposes $\mathbf{r}^{k}$ into smaller sub-sequences, denoted as $\mathbf{q}_{i}^{k}=\mathbf{r}^{k}\left[i*s:i*s+w\right]\quad 0\leq i\leq n_{s}=\frac{N^{k}}{s}$ which has a computation-efficiency length for the Transformer. The adjacent subsequences overlap by a distance of $s$, preventing the loss of local information, and $n_{s}$ is the number of subsequences.
  • We fed $\mathbf{q}_{i}^{k}$ into the TransBlock to learn the patterns inside the signal and get the features $\bar{\mathbf{q}}_{i}^{k}=\texttt{TransBlock}^{k}(\mathbf{q}_{i}^{k})$. After that, all the features $\bar{\mathbf{q}}_{i}^{k}$ of subsequences are gathered to form a new sequence $\textbf{t}=\left[\bar{\mathbf{q}}_{i}^{k}\right]_{i=0}^{n_{s}-1}$ and then passed into the next level.
Figure 4: Two voxels, denoted $v_{1}=(x_{1},y_{1},z_{1})$ and $v_{2}=(x_{2},y_{2},z_{2})$ that are located closely in the 3D space of MRI, but in the fMRI signals, they are far away from each other.
@@ 3.2.4 Brain Cognitive Features
Let $\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}}$ be the list of fMRI features that are the outputs from the Mutli-scale fMRI Transformer. $N_{r}$ is the number of regions in the brain. In particular, we define $N_{r}=6$ as discussed in the section 3.1. To learn the correlation between the ROIs, we adopt another feature embedding process by feeding $\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}}$ into a final TransBlock and receiving the brain cognitive features, denoted as $\mathbf{q}=\texttt{TransBlock}(\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}})$.
@@ 3.3 Training Objectives
Our framework is trained by a combination of Contrastive Loss, $\mathcal{L}_{con}$, and Brain fMRI Guidance Loss, $\mathcal{L}_{bfg}$, bellow:
$\small\begin{split}\mathcal{L}=\lambda_{con}\mathcal{L}_{con}+\lambda_{bfg}\mathcal{L}_{bfg}\end{split}$ (5)
where $\lambda_{con}$ and $\lambda_{bfg}$ are the weights of the loss functions. In this paper, we select $\lambda_{con}=\lambda_{bfg}=0.5$. The details of $\mathcal{L}_{con}$ and $\mathcal{L}_{bfg}$ are described in the sections below.
@@ 3.3.1 Contrastive Loss
$\small\begin{split}\mathcal{L}_{con}&=-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{p}_{i}\otimes\mathbf{q}_{i}/\sigma)}{\sum_{j}^{N}\exp(\mathbf{p}_{i}\otimes\mathbf{q}_{j}/\sigma)}\\ &-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{q}_{i}\otimes\mathbf{p}_{i}/\sigma)}{\sum_{j}^{N}\exp(\mathbf{q}_{i}\otimes\mathbf{p}_{j}/\sigma)}\end{split}$ (6)

### v2

# Brainformer: Mimic Human Visual Brain Functions to Machine Vision Models via fMRI

@@ Abstract
Human perception plays a vital role in forming beliefs and understanding reality. A deeper understanding of brain functionality will lead to the development of novel deep neural networks. In this work, we introduce a novel framework named Brainformer, a straightforward yet effective Transformer-based framework, to analyze Functional Magnetic Resonance Imaging (fMRI) patterns in the human perception system from a machine-learning perspective. Specifically, we present the Multi-scale fMRI Transformer to explore brain activity patterns through fMRI signals. This architecture includes a simple yet efficient module for high-dimensional fMRI signal encoding and incorporates a novel embedding technique called 3D Voxels Embedding. Secondly, drawing inspiration from the functionality of the brain’s Region of Interest, we introduce a novel loss function called Brain fMRI Guidance Loss. This loss function mimics brain activity patterns from these regions in the deep neural network using fMRI data. This work introduces a prospective approach to transfer knowledge from human perception to neural networks. Our experiments demonstrate that leveraging fMRI information allows the machine vision model to achieve results comparable to State-of-the-Art methods in various image recognition tasks.
@@ 1 Introduction
To avoid expensive annotation efforts, a surge in the development of foundational language models, e.g., BERT [14], GPT-2, GPT-3 [6], RoBERTa [44], T5[74], BART[39] is observed. These approaches are typically trained on extensive datasets and utilize the text as guidance for the vision models [93, 72, 29]. From the recent success of text as supervision for the visual learning, we revisit one of the key questions in the early days of artificial intelligence: What if human brain behaviors can serve as the guiding force for the machine vision models?. To answer this question, we found that Functional Magnetic Resonance Imaging (fMRI) has provided valuable insights into various aspects of human cognition and neuroscience. It contains rich information on how the human visual system works. For example, fMRI can help identify the specific regions of the brain that are active during various cognitive tasks. Studies [21, 22, 91]have shown that different intelligence-related tasks activate specific brain areas, such as problem-solving or memory. In addition, human intelligence is not just about the activity of individual brain regions but also about how different brain regions communicate. fMRI can reveal functional connectivity patterns, highlighting networks involved in various cognitive processes.
@@
While text typically reflects the outcomes of the recognition process, fMRI captures the dynamics of the cognitive processes involved. For instance, examiners are tasked with describing the context of an image, such as “A man is sitting on the chair." Initially, they focus on the man and then shift attention to the chair, determining the interactions between these objects, specifically, the act of sitting. Subsequently, they conclude the context in the text form. When encountering this description for the first time, it remains unclear which object the examiner attended to first or whether other smaller objects influenced their perception. Notably, this information can be encoded in fMRI signals. By analyzing specific brain regions or regions of interest, we can explore the whole recognition process of the human brain. This work provides a promising approach to further studies and endeavors to bridge the gap between human intelligence and deep neural networks. The contributions of this work can be summarized as follows.
Figure 1: Given a pair of images and fMRI signals (x-axis is the voxel index, y-axis is the magnitude of the voxel response), Brainformer can explore the local patterns of fMRI signals from brain regions and discover their interactions. Best view in color
@@
Contributions of this Work. This paper presents a new simple but efficient Transformer-based framework, named Brainformer, to analyze and leverage fMRI signals captured from human brain activities to supervise the machine vision learning model. Second, a new Brain fMRI Guidance Loss function is presented to distill information from fMRI extracted using Brainformer. It serves as a guidance signal to enhance the performance capabilities of vision models. Finally, Brainformer is designed for self-supervised learning and trained as an end-to-end deep network. It consistently outperforms previous State-of-the-Art Self-supervised learning methods across standard benchmarks, i.e., object detection, instance segmentation, semantic segmentation, and brain response prediction. The code will be released.
@@ 2 Related Work
Vision-Language Foundation Models. Pretraining on large-scale datasets [13, 50, 94] has become a popular approach for many visual recognition problems [27, 25, 59, 61, 56, 54, 55], e.g., classification, localization, or segmentation, etc. Despite this method has great success, it is limited by the number of annotated training data that is costly to collect. Recently, self-supervised training has been proposed to address this problem [60, 53, 58]. Significant advancements have occurred in vision-language pertaining [73, 29, 96, 66, 89, 87, 40, 92, 95, 49, 88]. This approach aims to encode both visual and text information into a model [82, 9, 97, 75, 36, 5, 53]. Recent work has proposed Image-Text foundation models that bridge the gap between image and text understanding [72, 30, 67, 89, 87]. They leverage dual-encoder architecture pre-trained on noise image-text pairs using contrastive learning or generative losses. Their studies showcase impressive performance in vision-language benchmarks while maintaining strong visual encoder capabilities for classification tasks.
Leveraging Human Brain Functions In Training. While the text has successfully demonstrated its efficiency in helping machine learning models enhance their performance, recent studies have been inspired by human brain mechanisms to improve models [84, 85, 32, 69, 41, 51, 18, 76, 43, 68, 34, 12, 77, 16, 38, 84, 71]. Some recent methods have used neural activity data to guide the training of models [65, 19, 80, 62, 52]. They utilized EEG and fMRI signals to constrain the neural network to behave like the neural response in the visual cortex.
@@
Decoding functional MRI. Decoding visual information from fMRI signals has been studied for a decade [24, 83, 31, 11, 23]. Most of these studies aimed to explore the hidden information inside the brain. It is a difficult task because of the low signal-to-noise ratio. Recently, with the help of deep learning, the authors in [10, 78, 81, 64, 41, 63] proposed methods to reconstruct what humans see from fMRI signal using diffusion models. Different from recent studies on decoding fMRI [81, 78, 10], our goal is to explore valuable vision information from fMRI signal, i.e., semantic and structure information, and utilize them as supervision for helping to enhance the recognition capability of the vision model. Although our approach also aims to extract feature representations of fMRI data, these previous methods still have limitations and do not apply to our problem. First, prior approaches, such as those by Yu et al. [81] and Paul et al., [78] relied on linear models and Multi-layer Perceptron to generate fMRI feature representations. This approach cannot capture the local pattern of the signals, especially the non-linear or complicated patterns. Kim et al. [35] presented SwiFT for self-training fMRI in 4D data, which does not apply to our problem. On the other hand, MindVis [10] utilized a Masked Brain Model inspired by MAE, which explored local patterns but ignored the correlations between multiple voxels, which describes how neurons interact inside the brain. Furthermore, MindVis failed to leverage the 3D spatial information in fMRI signals, which is crucial in understanding the relationships between voxels. Most notably, none of these approaches leverage characteristics of functional regions of interest (ROI) within the brain, which contain rich information about visual stimuli.
@@ 3 The Proposed Approach
3.1 Motivation
The fMRI contains a structured and semantically rich representation of information since fMRI captures brain activities in the context of a person engaging in visual tasks. It provides ground truth data about how the human brain responds to visual stimuli. According to recent studies [70, 33, 17], the brain can be divided into several Regions of Interest (ROI), where each region holds a different function. Especially in the scope of visual cognition, these are 6 regions specified as Early retinotopic visual regions (prf-visualrois), Body-selective regions (floc-bodies), Face-selective regions (floc-faces), Place-selective regions (floc-places), Word-selective regions (floc-words) and Anatomical streams (streams). In general, these regions are the responses for processing information related to the human body, face, place, motion, or objects regarding identification, recognition, etc. Therefore, the fMRI signals from these regions can be used to train computer vision models, providing a reference for what the brain is “seeing" during the training. By training vision models jointly with brain activities observed in fMRI scans, they can explore the relationship between the visual information processed by the brain and the visual information that models are learning. It can help uncover how the brain represents and processes visual information.
@@ 3.2 Brainformer
Figure 2: Brainformer utilizes fMRI signals (x-axis is the voxel index, y-axis is the magnitude of the voxel response) from specific brain regions as input, extracting the local features representing patterns within each region. The TransformerBlock measures the correlation among these regions to emulate brain activities. This information is subsequently transferred to the vision model through Contrastive Loss and Brain fMRI Guidance Loss.
3.2.1 High-dimensional fMRI Feature Encoding
@@
$\begin{split}\mathbf{r}^{k}&=\texttt{Conv1D}(m^{k})\in\mathbb{R}^{N^{k}\times d_{r}}\end{split}$ (2)
@@ 3.2.2 Brain 3D Voxel Embedding
$\begin{split}v_{i}^{k}&=\texttt{Linear}(x_{i}^{k},y_{i}^{k},z_{i}^{k})\in\mathbb{R}^{d_{r}}\\ v(m^{k})&=\texttt{concat}\left[v_{0}^{k},v_{1}^{k},\dots,v_{N^{k}-1}^{k}\right]\in\mathbb{R}^{N^{k}\times d_{r}}\end{split}$ (3)
@@ 3.2.3 Mutli-scale fMRI Transformer
Figure 4: Two voxels, denoted $v_{1}=(x_{1},y_{1},z_{1})$ and $v_{2}=(x_{2},y_{2},z_{2})$ that are located closely in the 3D space of MRI, but in the fMRI signals (x-axis is the voxel index, y-axis is the magnitude of the voxel response), they are far awya from each other.
@@
  • Step 1: We employ a slicing window with $w$ of width that traverses the entire signal $\mathbf{r}^{k}$ from its beginning to the end, with a step size of $s$. This process decomposes $\mathbf{r}^{k}$ into smaller sub-sequences, denoted as $\mathbf{q}_{i}^{k}=\mathbf{r}^{k}\left[i*s:i*s+w\right]\quad 0\leq i\leq n_{s}=\frac{N^{k}}{s}$ that has a computation-efficiency length for the Transformer. The adjacent subsequences overlap by a distance of $s$, preventing the loss of local information, and $n_{s}$ is the number of subsequences.
  • Step 2: We fed $\mathbf{q}_{i}^{k}$ into the TransBlock to learn the patterns inside the signal and get the features $\bar{\mathbf{q}}_{i}^{k}=\texttt{TransBlock}^{k}(\mathbf{q}_{i}^{k})$. After that, all the features $\bar{\mathbf{q}}_{i}^{k}$ of subsequences are gathered to form a new sequence $\textbf{t}=\left[\bar{\mathbf{q}}_{i}^{k}\right]_{i=0}^{n_{s}-1}$ and then passed into the next level.
@@ 3.2.4 Brain Cognitive Features
Let $\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}}$ be the list of fMRI features that are the outputs from the Mutli-scale fMRI Transformer. $N_{r}$ is the number of regions in the brain. In particular, we define $N_{r}=6$ as discussed in the section 3.1. To learn the correlation between the ROIs, we adopt another feature embedding process by feeding $\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}}$ into a final TransBlock and receiving the brain cognitive features, denoted as $\mathbf{q}=\texttt{TransBlock}(\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}})$
@@ 3.3 Training Objectives
Our framework is trained by a combination of Contrastive Loss and Brain fMRI Guidance Loss as $\mathcal{L}=\lambda_{con}\mathcal{L}_{con}+\lambda_{bfg}\mathcal{L}_{bfg}$ where $\lambda_{con}$ and $\lambda_{bfg}$ are the weights of the loss functions. In this paper, we select $\lambda_{con}=\lambda_{bfg}=0.5$. The details of $\mathcal{L}_{con}$ and $\mathcal{L}_{bfg}$ are described in the sections below.
@@ 3.3.1 Contrastive Loss
$\begin{split}\mathcal{L}_{con}&=-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{p}_{i}\otimes\mathbf{q}_{i}/\sigma)}{\sum_{j}^{N}\exp(\mathbf{p}_{i}\otimes\mathbf{q}_{j}/\sigma)}-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{q}_{i}\otimes\mathbf{p}_{i}/\sigma)}{\sum_{j}^{N}\exp(\mathbf{q}_{i}\otimes\mathbf{p}_{j}/\sigma)}\end{split}$ (5)